OpenStudy (anonymous):

Anyone?? Camp Pineveiw's cook, Margaret Johnson, was just about to begin preparing the picnic lunch for all the campers. She already knew she needed to fill 55 bowls of the same size and capacity with the same amount of food. When she was done, she decided to read the guidelines for the picnic, just out of curiosity. The guidelines said: 1. Every camper gets their own bowl of soup. 2. Every two campers will get one bowl of spaghetti to share. 3. Every three campers will get one bowl of salad to share. 4. All campers are required to have their own helping of salad, spaghetti, and soup.

OpenStudy (anonymous):

After some rapid calculations, Margaret was able to figure out how many campers were going to the picnic. Can you?

OpenStudy (anonymous):

@prowrestler @UnkleRhaukus

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@ybarrap

jimthompson5910 (jim_thompson5910):

This looks like a trial and error problem. We know that because "Every two campers will get one bowl of spaghetti to share" and because "Every three campers will get one bowl of salad to share." we can say that the number of campers is a multiple of 2*3 = 6

jimthompson5910 (jim_thompson5910):

So if there were 12 campers, then you'd have 12 bowls of soup 12/2 = 6 bowls of spaghetti 12/3 = 4 bowls of salad making 12+6+4 = 22 bowls total so that's too small

jimthompson5910 (jim_thompson5910):

If there were 18 campers, then you'd have 18 bowls of soup 18/2 = 9 bowls of spaghetti 18/3 = 6 bowls of salad making 18+9+6 = 33 bowls total still too small

jimthompson5910 (jim_thompson5910):

see how I'm doing this?

jimthompson5910 (jim_thompson5910):

actually, there's a much quicker way that I just thought about if x = # of campers, then x = # of bowls of soup x/2 = # of bowls of spaghetti x/3 = # of bowls of salad

OpenStudy (anonymous):

Try. you're on the right track.. Want a hint?

jimthompson5910 (jim_thompson5910):

Add them up and that total should be 55 x + x/2 + x/3 = 55 6x/6 + 3x/6 + 2x/6 = 55 I'll let you finish

OpenStudy (anonymous):

If there were only six campers, there would be six bowls of soup, three bowls of spaghetti, and two bowls of salad.

jimthompson5910 (jim_thompson5910):

I know. There are more than 6 campers.

jimthompson5910 (jim_thompson5910):

Solve that equation and you'll get the answer